Hybrid cars are touted as a "green" alternative; however, the financial aspects of hybrid ownership are not as clear. Consider the 2018 Edsel 550h, which had a list price of $5,000 (including tax consequences) more than the comparable Edsel 550. Additionally, the annual ownership costs (other than fuel) for the hybrid were expected to be $310 more than the traditional sedan. The EPA mileage estimate was 25 mpg for the hybrid and 17 mpg for the traditional sedan.

  

a.

Assume that gasoline costs $3.00 per gallon and you plan to keep either car for six years. How many miles per year would you need to drive to make the decision to buy the hybrid worthwhile, ignoring the time value of money? (Do not round intermediate calculations and round your answer to nearest whole number, e.g., 32.)

b. If you drive 14,000 miles per year and keep either car for six years, what price per gallon would make the decision to buy the hybrid worthwhile, ignoring the time value of money? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
c. Gasoline costs $3.00 per gallon and you plan to keep either car for six years. How many miles per year would you need to drive to make the decision to buy the hybrid worthwhile? Assume the appropriate interest rate is 10 percent and all cash flows occur at the end of the year. (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
d. If you drive 14,000 miles per year and keep either car for six years, what price per gallon would make the decision to buy the hybrid worthwhile? Assume the appropriate interest rate is 10 percent and all cash flows occur at the end of the year. (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

1.
=(5000/6+310)/(3*(1/17-1/25))=20246.5277777778

2.
=(5000/6+310)/(14000/17-14000/25)=4.33854166666667

3.
=(5000*10%/(1-1/1.1^6)+310)/(3*(1/17-1/25))=25819.4034696112

4.
=(5000*10%/(1-1/1.1^6)+310)/(14000/17-14000/25)=5.53272931491668